The only force acting on a 2.6 kg body as it moves along the positive x axis has an x component Fx =  −7x N, where x is in meters. The velocity of the body at x = 3.0 m is 8.2m/s.

The only force acting on a 2.6 kg body as it moves along the positive x axis has an x component Fx =-7x N, where x is in meters. The velocity of the body at x = 3.0 m is 8.2 m/s. (a) What is the velocity of the body at x 4.0 m? m/s (b) At what positive value of x will the body have a velocity of 5.0 m/s?
Answer #1

Her, from Newtons second law the expression for the force as follows: F = ma But the acceleration is the rate of change of v(a) Here, for x =4.0 m the velocity of the body becomes: +(1.3(x+(3.0m)) =(22+63 m/s -(1.35)|(4.0m) –(3.0m))= (13–168.2.199)